OpenStudy (anonymous):

Find critical numbers of the function f(x)=x^4-4x^2

OpenStudy (anonymous):

By the way this is a Calculus question.

OpenStudy (anonymous):

differentiate it and put that eqn equal to zero, the zeroes of that eqn is are the critical points

OpenStudy (anonymous):

i don't know what the answer for x^2-2=0 is

OpenStudy (anonymous):

take the first derivative of f(x) and you get: f'(x)=4x^3-8x then set that equal to zero 4x^3-8x=0 4x(x^2-2)=0 <--- Factored 4x=0 (x^2-2)=0 x=0 x^2=2 x= sqrt 2